OpenStudy (anonymous):

Factoring help? i have absolutely no idea how to do this, there are multiple problems like this and they didnt explain it very well. how would i go about solving this? Factor completely: xy + 4y − 2x − 8 (x − 4)(y + 2) (x + 4)(y − 2) (x − 4)(x − 2) Prime

OpenStudy (anonymous):

x(y-2) + 4(y-2) = (x+4)(y-2)

zepdrix (zepdrix):

Factoring is a little tricky if you've never done it before. You're pulling a value out of terms (dividing it out) but then leaving it as a multiplier on the outside of brackets. example: \[ab - 2a = a(b-2)\] see how we pulled an a out of each term, and left it as a multiplier on the outside? :O that is factoring. In your equation xy +y4 -2x -8, do you see any terms that share a variable that you can pull out of them? :D

OpenStudy (anonymous):

hmm, the 4 - 2?

OpenStudy (anonymous):

my apologies if im wrong, never done factoring before

zepdrix (zepdrix):

so a good place to start, is with the variables. notice that the first and 3rd term both have an x in them. so lets pull an x out of each. x(y-2) +4y - 8 understand what i did?

OpenStudy (anonymous):

oh, yeah, i kinda understand how you got that - so whats next?

zepdrix (zepdrix):

4y-8 , hmm they both seem to have a 4 in common. lets try pulling a 4 out of each term. 4(y-2)

OpenStudy (anonymous):

hmm, so we look for things that are similar or have a GCF then put that to the left?

zepdrix (zepdrix):

Yes it will almost always work out that way. There might be times when you don't want to pull out the GREATEST common term, you'll see why in a minute :U but it won't happen very often so the next step is kinda tricky, we've simplified the problem to x(y-2) + 4(y-2) yes?

OpenStudy (anonymous):

yes, after arranging them into that expression, what do we do now?

OpenStudy (anonymous):

btw, appreciate you explaining all this to me.

zepdrix (zepdrix):

So this is a little tricky. Notice that they both have a (y-2) term? so lets pull that out of each term :D and we get.. (y-2)[x+4]

OpenStudy (anonymous):

wow, factoring is weird. so we combine the two multipliers and the two things in the parentheses to get the answer then?

zepdrix (zepdrix):

you are pulling OUT the parentheseed(is that a word..? XD) term from each .. and then looking at what is left over x(y-2) + 4(y-2) = (y-2) (x*_____ + 4*_____ ) since we divided out the (y-2), the space is just a 1 left over there on each term :O hopefully i didn't make that too confusing..

OpenStudy (anonymous):

ohh, alright :) well ty for helping, really appreciate it